A GP consist of even number of terms. The sum of all the terms is 3 times that of odd terms, find common ratio. 

Grade:12th pass

1 Answers

Susmita
5 years ago
Let us consider a series which has 4 terms,viz,a1,a2,a3,a4.
According to the question
a1+a2+a3+a4=3(a1+a3)
Or,-2a1+a2-2a3+a4=0
Let r be the common ratio.
Then a2=a1r,a3=a1r2,a4=a1r3.Put it in the equation.a1 cannot be zero.so
-2+r-2r2+r3=0
Or,r2(r-2)+1(r-2)=0
Or,(r-2)(r2+1)=0
either r=2 or r2=-1(r would be imaginary in this case,so you can exclude it).
r=2